OpenStudy (anonymous):

1/4*3/5 (those are fractions not division) Will medal and fan

OpenStudy (anonymous):

@KyleWells17 can you help me?

OpenStudy (sneakysquash):

i could help dude/

OpenStudy (sneakysquash):

so its 1/4 times 3/5 right?

OpenStudy (anonymous):

yea

OpenStudy (sneakysquash):

ight.

OpenStudy (sneakysquash):

???

OpenStudy (sneakysquash):

sorry dude, nm.

OpenStudy (sneakysquash):

idk, i always get lost when i comes to making the denominators the same.

OpenStudy (anonymous):

Its ok :)

OpenStudy (benlindquist):

@ash_heather1 what is the gcf

OpenStudy (anonymous):

20?

OpenStudy (benlindquist):

20

OpenStudy (benlindquist):

1 multiplied by 3 is

OpenStudy (anonymous):

yea then do to the top what you did to the bottom right?

OpenStudy (benlindquist):

i think so

OpenStudy (benlindquist):

what is ur answer

OpenStudy (anonymous):

5/20+12/20=17/20

OpenStudy (benlindquist):

wheredid the 5 and 12 come from

OpenStudy (anonymous):

since 4*5=20 i have to do 1*5 and since i did 5*4=20 i had to do 2*4

OpenStudy (benlindquist):

ok, i believe you are correct

563blackghost (563blackghost):

incorrect... a simplier way to find the answer is too multiply across...\[\frac{ 1 }{ 4 } \times \frac{ 3 }{ 5 } = \frac{ ? }{ 20 }\] what i typed eariler was incorrect plz forgive me

OpenStudy (benlindquist):

Cross multiply, so 1x5 and 4x3 = what /20

OpenStudy (anonymous):

isint it 17?

OpenStudy (benlindquist):

yes

OpenStudy (benlindquist):

you final answer

OpenStudy (anonymous):

17/20

563blackghost (563blackghost):

hmmm thats not what i got i got 3/20...

OpenStudy (benlindquist):

ok

OpenStudy (benlindquist):

3/20 what

OpenStudy (benlindquist):

i thought you cross multiply

563blackghost (563blackghost):

we are multiplying the fraction not cross multiplying... you use cross multiplication to find out which fraction is larger but not when multiplying two fracions into one

OpenStudy (benlindquist):

3/20

OpenStudy (anonymous):

ok im confused...

563blackghost (563blackghost):

im srry.... example of cross multiplication.. https://www.mathsisfun.com/algebra/cross-multiply.html when doing problems like yours you would simply multiply across so 1 x 3 and 4 x 5

OpenStudy (anonymous):

oh ok i got it! thanks

563blackghost (563blackghost):

np glad u got it ^^